Light striking a metal surface causes electrons to be emitted from the metal via the photoelectric effect.
In a particular experiment to study the photoelectric effect, the frequency of the incident light and the temperature of the metal are held constant. Assuming that the light incident on the metal surface causes electrons to be ejected from the metal, what happens if the intensity of the incident light is increased?
Check all that apply.
The work function of the metal decreases.
The number of electrons emitted from the metal per second increases.
The maximum speed of the emitted electrons increases.
The stopping potential increases.
Answer #1
Concept and reason

The concept needed to solve this problem is the photo electric effect.

Explain about the technical terms work function and stopping potential in photo electric effect. After that, explain how the speed of the emitted electrons increases in the photo electric effect. Finally, explain what happens if the intensity of the photo light is increased.

Fundamentals

Photo electric effect is a physical phenomenon in which electrons are removed from the photo metals by using suitable light.

Work function:

Work function is the characteristic of the photo metal which represents the minimum energy of the light is needed to start the photo electric effect.

In the photo electric effect, the sum part of the energy of the incident photon on the metal is useful to start the photo electric effect that is equal to work function (ϕ)\left( \phi \right) of the metal and the other part of the energy of the incident light is useful to increase the kinetic energy (K) of the emitted electrons.

E=ϕ+KE = \phi + K

Here, E is the total energy of the incident light.

Stopping potential:

The potential difference needed to stop the emitted electrons. If the speed of the emitted electrons is increased, the stopping potential increases to stop the electrons and vice versa.

If the frequency of the incident light is increased, then the energy of the incident light increases. So, the value of kinetic energy of the emitted electrons will be increased as per the equation E=ϕ+KE = \phi + K . Here, the value of ϕ\phi is constant for a given photo metal.

If the intensity of a light is increased, then the number of photons in the light increases. As one photon is responsible to remove one electron from the metal, the number of emitted electrons increases if the intensity of the light is increased.

From the definition of work function, the option the work function of the metal decreases is incorrect.

From the definition of stopping potential, the option the stopping potential increases is incorrect.

The value of kinetic energy of the emitted electrons will be increased as per the equation E=ϕ+KE = \phi + K . Here, the value of ϕ\phi is constant for a given photo metal. So, the option the maximum speed of the emitted electrons increases is incorrect.

If the intensity of a light is increased, then the number of photons in the light increases. As one photon is responsible to remove one electron from the metal, the number of emitted electrons increases if the intensity of the light is increased.

Hence, the option the number of electrons emitted from the metal per second increases is correct.

Ans:

The option the number of electrons emitted from the metal per second increases is correct.